OpenStudy (anonymous):

Find the least common denominator (LCD) for the fractions. 1/6 3/10

OpenStudy (anonymous):

30 because 6 = 2 * 3 and 10 = 2 * 5 so they have 2 in common. so now we multiply 2 * 3 * 5 to get 30.

OpenStudy (anonymous):

LCD is the lowest amount of times two or more numbers can go into another. 6 and 10 both go into 30 evenly. (6 goes into 30 5 times and 10 goes into 30 3 times) 1/6 would be changed to 5/30. 3/10 would be changed to 9/30. The LCD is 30.

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

could u do a couple more

OpenStudy (anonymous):

i mean help me with some more

OpenStudy (anonymous):

Choose the missing number to make equivalent fractions. 5/6 = 20/?

OpenStudy (anonymous):

5 * 4 = 20 6 * 4 = ?

OpenStudy (anonymous):

i have to choose the missing number the options are 21 22 24 and 30

OpenStudy (anonymous):

what is 6 times 4, that's your answer.

OpenStudy (anonymous):

I'll give you a hint, it's C. 24

OpenStudy (anonymous):

oh ok i got u thxs

OpenStudy (anonymous):

Which number can replace the variable to make these fractions equivalent? n/7 = 45/52

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

4 8 6 9

OpenStudy (anonymous):

I'd go with 6

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Which is a correct comparison of the fractions? 7/12 and 11/20 A 7/12 <11/20 B 7/12>11/20 C 7/12=11/20

OpenStudy (anonymous):

what do u think

OpenStudy (anonymous):

B. 7/12 is definitely greater then 11/20